# Multilingual hierarchical classification of job advertisements for job vacancy statistics

### v1

@@ 7.1 Results for country specific classifier (trained on Polish language only)
Figure 5 compares the accuracy of different classification approaches for job codes across various levels of the KZIS occupation classification system, ranging from 1-digit to 6-digit codes, with ISCO classification at the 4-digit level. Across all datasets (Overall, CBOP, and Hand-coded job offers), transformer-based models consistently outperform linear models. Larger transformer models (HerBERT-large and XLM-RoBERTa-large) achieve the highest accuracy scores across all levels and datasets but are slightly better than base models.
Figure 5: Results for the recall@1 classification of Polish job advertisements by classifier
Notably, the top-down approach consistently yields better results than the bottom-up approach for all models and coding levels. This improvement is particularly pronounced for hand-coded data, indicating that taking the hierarchical structure of occupational codes into account significantly enhances prediction accuracy. For instance, for the hand-coded dataset for HerBERT-base the difference at ISCO level is close to 1.5p.p. The same relation is present for the linear model where the differences between top-down and bottom-up are close to 2 p.p. (e.g., ISCO level)
As expected, the accuracy of all models decreases as the level of detail in the occupational codes increases. For example, the best-performing model (XLM-RoBERTa-large and HerBERT-large with top-down approach) achieves around 88.6% accuracy for 1-digit codes but drops to about 72.3% for 6-digit codes. This trend is consistent across all models and approaches, reflecting the increasing difficulty of classification as the categories become more specific.
@@ 7.2.1 Overall results
Figure 6 compares the accuracy (measured as Recall@1) of different classifiers for occupational coding across various levels of detail, from 1-digit to 6-digit codes, as well as ISCO codes. The classifiers include different transformer models (HerBERT and XLM-RoBERTa) using both top-down and bottom-up approaches. The plot also compares these automated methods to hand-coded results.
@@ 7.2.2 Detailed results for the hand-coded data
This section presents a more detailed analysis of the combined hand-coded data set, comprising 10,000 and 1,000 records. The data primarily come from online sources, and the classifier will be employed for further analysis of this type of data. Table 11 contains the corresponding data for the transformer model. The results are presented separately for the training and test data. For the training dataset, the transformer model achieves over 92% agreement for each group, while the logistic model performs significantly worse for the 1, 3, 4, and 9 groups.
@@ 7.3 Results for multilingual dataset (24 languages)
In this section, we present the results for a multilingual dataset that has been trained with XLM-RoBERTa models. Table 12 presents the results for the test dataset only. The table also includes two approaches to modelling the hierarchy for 1, 2, 4 (ISCO) and 6 (KZiS) digits. The results are comparable to those observed for the Polish and bilingual models. The overall classification rate for 1-digit codes is approximately 84% (approximately 2-4 p.p. lower than the first two approaches), and the hand-coded classification rate is approximately 78%. As previously noted, the inclusion of the hierarchy improves predictions for the hand-coded data as the hierarchy is traversed.
@@
As indicated in Section 5.2, the sample size for languages differs considerably in the test data. Accordingly, in order to evaluate the accuracy of the multilingual classifier, we present the results for each translated test data set, including the complete data set and the sample utilized for training. Figure 7 depicts the accuracy (recall@1) for each language across 1-digit, 4-digit (ISCO) and 6-digit (KZiS) occupation codes, employing both the bottom-up and top-down approaches. The figures illustrate the results for the complete and sampled hand-coded test data sets.
@@
A modest correlation was observed between the accuracy of the XLM-Roberta model and the coverage of languages. This is evidenced by the Pearson’s and Spearman’s correlation coefficients between accuracy and the share of tokens, which were 0.26 and 0.17, respectively (see Table 10). The highest Pearson’s and Spearman’s correlation coefficients were observed for the share of cases in the test data (0.37 and 0.22, respectively). The detailed results for this classifier are presented in the Supplementary Materials.
@@ Acknowledgements
Cherniaiev Herman, Pater Robert, Wydmuch Marek work and research were conducted in the Educational Research Institute (IBE) in Warsaw, and funded by the Polish Ministry of Education and Science within a project Supporting IQS functioning and improvements in order to use its solutions in achieving the country’s development strategy aims, ZSK5 (POWR.02.13.00-IP.02-00-002/20). The preparation of hand-coded datasets was financed by IBE. Maciej Beręsewicz was partially supported by the Polish National Agency for Academic Exchange (NAWA) under the project number BPN/BEK/2023/1/00099.
@@
    Additional information on the data and its quality (in PDF)
@@ 8 CBOP 61,596
cloud||||cloud computing||||chmura 24 4
@@
uczenie maszynowe||||machine learning||||data scientist 21 5
@@
informatyk(a||||i) przemysłow(a||||ej)||||industrial computer science 11 4
@@
blockchain||||block chain 9 3
@@
internet rzeczy||||IoT||||internet of things 6 2
@@
(frontend||||front end) developer 4 1
@@
pianino||||fortepian||||pianina 4 2
virtual reality||||wirtualna rzeczywostość 3 2
@@
Note: Table 17 provides information on the number of unique 6-digit and 1-digit codes for a given stratum defined by phrase. This can provide information about the quality of the initial stratification based on keyword searches. For example, the main codes for ’pianino|fortepian|pianina’ were related to kitchen and restaurant organisation. The reason for this is that these keywords are related to the name of the restaurant rather than the pianino itself.

### v2

@@ 7.1 Results for country specific classifier (trained on Polish language only)
Figures 5-7 compares the accuracy of different classification approaches for job codes across various levels of the KZIS occupation classification system, ranging from 1-digit to 6-digit codes, with ISCO classification at the 4-digit level. Across all datasets (Overall, ePraca, and Hand-coded job offers), transformer-based models consistently outperform linear models. Larger transformer models (HerBERT-large and XLM-RoBERTa-large) achieve the highest accuracy scores across all levels and datasets but are slightly better than base models.
Figure 5: Results for the recall@1 classification of Polish job advertisements by classifier (Overall)
Figure 6: Results for the recall@1 classification of Polish job advertisements by classifier (ePraca only)
Figure 7: Results for the recall@1 classification of Polish job advertisements by classifier (Hand-coded only)
Notably, the top-down approach consistently yields better results than the bottom-up approach for all models and coding levels. This improvement is particularly pronounced for hand-coded data, indicating that taking the hierarchical structure of occupational codes into account during model training significantly enhances prediction accuracy. For instance, for the hand-coded dataset for HerBERT-base the difference at ISCO level is close to 1.5p.p. The same relation is present for the linear model where the differences between top-down and bottom-up are close to 2 p.p. (e.g., ISCO level)
As expected, the accuracy of all models decreases as the level of detail in the occupational codes increases. For example, the best-performing model (XLM-RoBERTa-large and HerBERT-large with top-down approach) achieves around 88.6% accuracy for 1-digit codes but drops to about 61% for 6-digit codes. This trend is consistent across all models and approaches, reflecting the increasing difficulty of classification as the categories become more specific.
@@ 7.2.1 Overall results
Figure˜8 compares the accuracy (measured as Recall@1) of different classifiers for occupational coding across various levels of detail, from 1-digit to 6-digit codes, as well as ISCO codes. The classifiers include different transformer models (HerBERT and XLM-RoBERTa) using both top-down and bottom-up approaches. The plot also compares these automated methods to hand-coded results.
@@ 7.2.2 Detailed results for the hand-coded data
This section presents a more detailed analysis of the combined hand-coded data set, comprising 10,000 and 1,000 records. The data primarily come from online sources, and the classifier will be employed for further analysis of this type of data. Table˜11 contains the corresponding data for the transformer model. The results are presented separately for the training and test data. For the training dataset, the transformer model achieves over 92% agreement for each group, while the logistic model performs significantly worse for the 1, 3, 4, and 9 groups.
@@ 7.3 Results for multilingual dataset (24 languages)
In this section, we present the results for a multilingual dataset that has been trained with XLM-RoBERTa models. Table˜12 presents the results for the test dataset only. The table also includes two approaches to modelling the hierarchy for 1, 2, 4 (ISCO) and 6 (KZiS) digits. The results are comparable to those observed for the Polish and bilingual models. The overall classification rate for 1-digit codes is approximately 84% (approximately 2-4 p.p. lower than the first two approaches), and the hand-coded classification rate is approximately 78%. As previously noted, the inclusion of the hierarchy improves predictions for the hand-coded data as the hierarchy is traversed.
@@
As indicated in Section˜5.2, the sample size for languages differs considerably in the test data. Accordingly, in order to evaluate the accuracy of the multilingual classifier, we present the results for each translated test data set, including the complete data set and the sample utilized for training. Figure˜9 depicts the accuracy (recall@1) for each language across 1-digit, 4-digit (ISCO) and 6-digit (KZiS) occupation codes, employing both the bottom-up and top-down approaches. The figures illustrate the results for the complete and sampled hand-coded test data sets.
@@
A modest correlation was observed between the accuracy of the XLM-Roberta model and the coverage of languages. This is evidenced by the Pearson’s and Spearman’s correlation coefficients between accuracy and the share of tokens, which were 0.26 and 0.17, respectively. The highest Pearson’s and Spearman’s correlation coefficients were observed for the share of cases in the test data (0.37 and 0.22, respectively). The detailed results for this classifier are presented in the Supplementary Materials.
@@ 8 Discussion
The paper highlighted the importance of data quality as an input to the training and testing models. We have shown that the hand-coded and administrative data can introduce additional uncertainty due to misclassification of labels and hierarchical classification. This process should certainly be included in the National Statistical Quality Assurance Frameworks and the European Statistics Code of Practice in order to progressively include indicators on the quality of training and testing data used for official statistics.
@@ Acknowledgements
Cherniaiev Herman, Pater Robert, Wydmuch Marek work and research were conducted in the Educational Research Institute – National Research Institute (IBE-PIB) in Warsaw, and funded by the Polish Ministry of Education and Science within a project Supporting IQS functioning and improvements in order to use its solutions in achieving the country’s development strategy aims, ZSK5 (POWR.02.13.00-IP.02-00-002/20). The preparation of hand-coded datasets was financed by IBE. Maciej Beręsewicz was partially supported by the Polish National Agency for Academic Exchange (NAWA) under the project number BPN/BEK/2023/1/00099. In addition Cherniaiev Herman, Pater Robert and Maciej Beręsewicz work supported by National Science Centre, Poland within a project OJALAB: Online job advertisements to study skill demand and job search patterns [Grant number: 2024/53/B/HS4/01580].
@@
    Additional information on the data and its quality (Section A).
@@ Appendix A Additional information on the data and its quality
cloud||cloud computing||chmura 24 4
@@
uczenie maszynowe||machine learning||data scientist 21 5
@@
informatyk(a||i) przemysłow(a||ej)||industrial computer science 11 4
@@
blockchain||block chain 9 3
@@
internet rzeczy||IoT||internet of things 6 2
@@
(frontend||front end) developer 4 1
@@
pianino||fortepian||pianina 4 2
virtual reality||wirtualna rzeczywostość 3 2
@@
Note: Table˜17 provides information on the number of unique 6-digit and 1-digit codes for a given stratum defined by phrase. This can provide information about the quality of the initial stratification based on keyword searches. For example, the main codes for ’pianino|fortepian|pianina’ were related to kitchen and restaurant organisation. The reason for this is that these keywords are related to the name of the restaurant rather than the pianino itself.